OpenStudy (mony01):

verify that the given function satisfies both the differential equation and the initial condition. y'+3y=e^-t y(0)=-1/2 y=((e^-t)/2)-e^(-3t)

OpenStudy (tkhunny):

You should NEVER stumble over a "verify". Just do it. You have y. Find y'. Check it out.

OpenStudy (mony01):

I don't know how to do the derivative of y

OpenStudy (tkhunny):

Find these derivatives: \(y = e^{t}\) \(y = e^{-t}\) \(y = ½e^{-t}\) \(y = e^{-3t}\) Please demonstrate proficiency.

OpenStudy (mony01):

is the derivative for y=e^t is y'=e^t y=e^-t is y'=-e^-t y=e^-3t is y'=-3e^-3t

OpenStudy (tkhunny):

The third one seems to be missing.

OpenStudy (mony01):

i cannot see what you wrote is it the same as y=e^-t

OpenStudy (tkhunny):

Oh, right. Silly coding issues. \(y = \dfrac{1}{2}e^{-t}\)

OpenStudy (mony01):

\[y'=-\frac{ e ^{-t} }{ 2 }\]

OpenStudy (tkhunny):

That's the spirit. Guess what? There is now nothing in your way that can prevent you from finding that derivative in the original problem statement.

OpenStudy (mony01):

is it \[y'=-\frac{ e ^{-t} }{ 2}+3e ^{-3t}\]

OpenStudy (tkhunny):

There it is! Okay, substitute and go!

OpenStudy (mony01):

what do i do with the y(0)?